OpenStudy (anonymous):

How does the volume of a cone change when the radius is quadrupled and the height id reduced to 1/5 of the original size?

OpenStudy (anonymous):

Volume of cone = V = 1/3 (pi) (r^2) (h) New cone = V2 = 1/3 (pi) (4^2 r^2) (1/5 h) Compare V with V2 and factorize out 1/3 (pi) (r^2) (h) for V2